OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of two thirds?

OpenStudy (anonymous):

2x + 3y = 3 2x − 3y = 3 3x + 2y = 3 3x − 2y = 3

OpenStudy (anonymous):

2x − 3y = 3